OpenStudy (anonymous):

What is the ratio of x to y? 3/2=2x/2x+y So this means x=1.5 and y is equal to -1, but how does that help me, when my options are: A. -3/+2 B.-2/+3 C. 2/3 D. 3/2 I was looking it up and I believe it is a, I would like clarification though, please help!

OpenStudy (anonymous):

@pooja195 or anyone else who can help me.

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (anonymous):

Will fan and mettle please help!

rebeccaxhawaii (rebeccaxhawaii):

im sorry i forgot this already but ill summon @AloneS

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Hey guys, can you help me? (a wolf lover and Alone S

OpenStudy (anonymous):

)*

OpenStudy (mrnood):

Your initial statement is wrong you cannot get a value for x AND y from a single equation

rebeccaxhawaii (rebeccaxhawaii):

omg mrmood you're almost 90 SS

OpenStudy (anonymous):

how come? 2x is equal to three..... Meaning 2*1.5=3... So x is equal to 1.5 and same for y. Please explain

OpenStudy (mrnood):

\[\frac{ 3 }{ 2 } = \frac{ 2x }{2x+y }\] first 'cross multiply' to get rid of the fractions

OpenStudy (anonymous):

So .3*2x+y and 2*2x

OpenStudy (mrnood):

3(2x+y) = 2(2x) expand that then arrange to get y terms on one side, and x on the other

OpenStudy (anonymous):

3+y=2(4x)

OpenStudy (mrnood):

no expand the brackets (distribute the 3 on left side and the 2 on right side)

OpenStudy (mrnood):

Don't rush or guess

OpenStudy (anonymous):

Can you show me? I'm not sure exactly what you are saying to do...

OpenStudy (mrnood):

the left side of the equation is 3(2x+y) that is the same as 3 times (2x +y) so write that in full

OpenStudy (anonymous):

What do you mean though "in full"?

OpenStudy (mrnood):

 multiply all terms in the bracket by 3